A 5.0g conducting sphere with charge of 20 μC hangs by a non-conducting thread in an electric field produced by two plates separated by 8.0cm. What potential will cause the ball to hang at 25° to the vertical? 


1
Expert's answer
2021-01-19T07:12:05-0500

Let’s write the conditions of the equilibrium for the conducting sphere:


"\sum F_x=0, \sum F_y=0"

Let’s consider the forces that act on the conducting sphere in the horizontal "x"- and vertical "y"-direction:


"Tsin\theta-F_e=0, (1)""Tcos\theta-mg=0. (2)"

We can express the force "T" from the second equation:


"T=\dfrac{mg}{cos\theta}."

Then we can substitute it into the first equation and find the electrostatic force, "F_e":


"mgtan\theta-F_e=0,""F_e=mgtan\theta."

From the other hand, "F_e=qE." From this formula we can find the electic field strength:


"E=\dfrac{F_e}{q}=\dfrac{mgtan\theta}{q}."

Finally, we can find the potential that will cause the ball to hang at 25° to the vertical from the formula:


"\Delta V=Ed=\dfrac{mgdtan\theta}{q},""\Delta V=\dfrac{5.0\cdot10^{-3}\ kg\cdot 9.8\ \dfrac{m}{s^2}\cdot 8\cdot10^{-2}\ m\cdot tan25^{\circ}}{2\cdot10^{-5}\ C}=91.4\ V"

Answer:

"\Delta V=91.4\ V"